OpenStudy (anonymous):

how to integrate 2/(x^2)+1) thx

OpenStudy (dumbcow):

trig substitution x = tan(theta) dx = sec^2(theta) --> tan^2 + 1 = sec^2

OpenStudy (anonymous):

oh,get it thx

OpenStudy (dumbcow):

:)